OpenStudy (anonymous):

Tim is twice as old as his son. In six years, Tim's age will be three times what his son's age was six years ago. How old is Tim's son now?

OpenStudy (anonymous):

T = 2S T-6 = 3(S-6) substitute the first into the second equation: 2S - 6 = 3S - 18 -6 = S - 18 12 = S so S=12, T=24.... Tim is 24 years old now and his son is 12. If Tim is a gun, then his son is a son of a gun.... :)

OpenStudy (anonymous):

LOLOL!xD